OpenStudy (anonymous):

2 square root 32 + 3 square root 72 HELP!

OpenStudy (anonymous):

\[2\sqrt{32}+3\sqrt{72}=2\sqrt{16*2}+3\sqrt{36*2}\] \[8\sqrt{2}+18\sqrt{2}=26\sqrt{2}\]